OpenStudy (anonymous):

A homeowner plots her house and yard on a computer. She determines that the circle that defines the part of the front yard that gets watered by the sprinkler is (x – 9)2 + (y + 11)2 = 16. What is the diameter (in meters) of the circular area that gets watered by the sprinkler? 4m 8m 16m 32m its either of those answers! thanks so much!

OpenStudy (michele_laino):

hint: the equation of a circumference, whose center is the point C=(a,b) and whose radius is R, is given by the subsequent formula: \[{\left( {x - a} \right)^2} + {\left( {y - b} \right)^2} = {R^2}\]

OpenStudy (michele_laino):

furthermore, I can rewrite your equation as follows: \[{\left( {x - 9} \right)^2} + {\left( {y - \left( { - 11} \right)} \right)^2} = {4^2}\]

OpenStudy (michele_laino):

So, what is the radius of your circular area?